# Semantic Exploration and Dense Mapping of Complex Environments using Ground Robots Equipped with LiDAR and Panoramic Camera

### v1

@@ V-D Physical Experiments
In the 12 m × 8 m mock construction site with 5 semantic targets (3 window frames, 1 work platform, and 1 cone), we used YOLOv7 pre-trained on construction data [26] for object detection and SAM2[27] for segmentation. The robot navigated at an average speed of 0.7 m/s (linear) and 0.4 rad/s (angular), completing the exploration in 54 seconds. Fig. 5(b) shows the trajectory and the overall point cloud, including the dense background and reconstructed semantic objects. Figs. 5(c)-(j) present the comparisons between reconstructed and ground-truth objects. The system achieved 87.19% object map completeness (points within 2.5 cm of the ground truth) with an average error of 3.11 cm, and 99.19% background map completeness with a 2.27 cm error.
In the 12 m × 23 m lobby with 11 chairs of diverse types as semantic targets, the exploration was completed in 132 seconds at 0.9 m/s (linear) and 0.6 rad/s (angular). the robot followed a consistent trajectory that efficiently covered the environment and ensured the observation of all the viewpoints of the targets (Fig. 1). Fig. 6 presents the reconstructed objects and the error analysis. The object map achieved 97.28% completeness with an average error of 1.75 cm, while the background map reached 99.33% completeness and 2.31 cm average error.
Figure 6: Mapping results and error visualization for the lobby scene. (a)-(b) Ground-truth point clouds of two example chairs. (c)-(d) Reconstructed point cloud with error heatmaps. (e)Scene-wide error heatmap comparing reconstruction to ground truth. The error increases from blue, white to red.
VI CONCLUSIONS
This paper presents a comprehensive system for autonomous semantic exploration and dense semantic target mapping. To address the challenge of balancing observation quality and exploration efficiency, we propose a hierarchical planning architecture that handles semantic and geometric viewpoints explicitly while considering their mutual impact, a safe-aggressive exploration state machine that supports aggressive strategy, and a modular mapping component that achieves accurate and complete semantic target mapping utilizing off-the-shelf SLAM algorithms. Extensive simulation experiments validate the planning efficiency, while two real-world deployments demonstrate the system’s ability to achieve accurate and complete semantic mapping with minimal exploration overhead. Future work could extend the system to heterogeneous-robots collaboration.
ACKNOWLEDGMENT
This work was supported in part by YKK AP Inc. The authors gratefully acknowledge their support.

### v2

@@ V-D Physical Experiments
In the 12 m × 8 m mock construction site with five semantic targets (three window frames, one work platform, and one cone), we used YOLOv7 pre-trained on construction data [31] for object detection and SAM2 for segmentation. The robot navigated at an average speed of 0.7 m/s (linear) and 0.4 rad/s (angular), completing the exploration in 54 seconds. Fig. 5 provides the visualization of the exploration trajectory, complete map, and error heatmap. The system achieved 87.19% object map completeness (points within 2.5 cm of the ground truth) with an average error of 3.11 cm, and 99.19% background map completeness with a 2.27 cm error. In the 12 m × 23 m lobby with 11 chairs of diverse types as semantic targets, we used YOLOV8 instance segmentation pretrained on COCO dataset [28] to provide object detection and segmentation. The exploration was completed in 132 seconds at 0.9 m/s (linear) and 0.6 rad/s (angular). As illustrated in Fig. 1, the robot followed a consistent trajectory that efficiently covered the environment and ensured the observation of all the viewpoints of the targets. Fig. 6 presents the reconstructed objects and the error analysis. The object map achieved 97.28% completeness with an average error of 1.75 cm, while the background map reached 99.33% completeness and 2.31 cm average error. On average, the memory cost per occupied voxel to store a dense cloud is 4,179.70 bytes, and the total memory cost increases linearly with time. Given the 64 GB RAM of our device, there is ample headroom for a larger-scale environment.
Figure 6: Map and error visualization for the lobby scene. (a)-(b) Ground-truth point clouds of two chairs. (c)-(d) Reconstructed point cloud with error heatmaps. (e) Scene-wide error heatmap comparing. The error increases from blue to red.
VI Conclusion
This paper presents a complete system for autonomous semantic exploration and dense target mapping. We propose a hierarchical planner that explicitly handles semantic and geometric viewpoints, a safe-aggressive exploration state machine for efficient coverage, and a modular mapping component that leverages off-the-shelf SLAM for accurate semantic mapping. Extensive simulations validate planning efficiency, while real-world deployments demonstrate precise and complete mapping with minimal exploration effort.